OpenStudy (anonymous):

Solve for x. Enter only a number. 3 = log8 + 3logx.

OpenStudy (mertsj):

\[3=\log8 + logx^3\] \[3=\log(8x^3)\] \[10^3=8x^3\] \[1000=8x^3\] \[125=x^3\] \[\sqrt[3]125{}=\sqrt[3]{x^3}\] \[5=x\]